The volume (in mL) will a sampleof F_2 gas occupy
Answer #1

from P1V1= P2V2

V2 =25 ml

P2=1.2 atm

P1= 5.5 atm

V1= P2V2/P1= 25*1.2/5.5 =5.45 ml. Correct answer is b.